With the weather over the last week and especially this weekend being sunny there were great expectations of change since the last apiary inspection.
Hive 2 had some dead bees outside the hive and on the hive floor, stores were still low and the bee population although over 9 frames was low on each frame. There were a couple of bees with deformed wing on the mesh floor but those on the comb seemed fine, plus there were significant amount of chalk brood, indicating chilled brood which would stack up with the dispersed bees across the brood comb. The comb needs changing in this hive and the original plan was to carry out an artificial swarm once the colony had built up in order to make increase, may need to reconsider the plan. Left feed on this hive, cleared dead bees from in front of hive, changed floor and will monitor to see if things change.
The other two hives were a better story. We are carrying out a bailey comb change on hive 5 and the bees are beginning to draw out the comb in the new brood chamber. Hive 7 now has sufficient stores and number of bees have increased so we will continue with the plane to artificial swarm this colony to make increase sometime in the near future.